How would I make 1/6^ (up something) = 6? Thank you much if you can help.
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

How would I make 1/6^ (up something) = 6?
--------------------------
Ans: (1/6)^(-1) = 6
======================
Cheers,
Stan H.
----------------------

Answer by richwmiller(17219)


(1/6)^-1 =6
or
1/(6^-1)=6 1/(1/6)